OpenStudy (anonymous):

0.2f + 0.6(f+20)= -8+0.4f Identity or no solution and how ?

OpenStudy (anonymous):

0.2f + 0.6f + 12 = -8 + 0.4f (distributive property used) 0.8f + 20 = 0.4f 0.4f = -20 f = -50 (the one solution) The original equation was neither an identity nor a "no solution" equation.